# A minimal model with stochastically broken reciprocity

### v1

@@ Abstract
We introduce a minimal model consisting of a two-body system with stochastically broken reciprocity (i.e., random violation of Newton’s third law) and then investigate its statistical behaviors, including correlation functions, time evolution of probability distribution functions, energy gain, and entropy production. The effective temperature of this two-body system immersed in a thermal bath is derived. Furthermore, we heuristically present an extremely minimal model where the relative motion adheres to the same rules as in classical mechanics, while the effect of stochastically broken reciprocity only manifests in the fluctuating motion of the center of mass.
Keywords: stochastically broken reciprocity, correlation function, probability distribution function, energy gain, entropy production
@@ I Introduction
The strict breaking of reciprocity leads to odd viscosity or elasticity [25, 26, 29, 27, 28, 31, 30], unconventional phase transitions [32, 33, 34, 35, 36, 37, 38], and exotic transport behaviors [39, 40, 41, 42]. Rather than focusing on the aforementioned in-depth discussions on consequences of a strict violation of reciprocity, we aim to explore a more delicate scenario: Newton’s third law holds on average, but reciprocity is broken stochastically. Since the third law holds in average, we expect that the core results of classical mechanics remain valid when expressed in terms of the mean values of relevant physical quantities. Stochastically broken reciprocity, however, plays a role in the fluctuations of these physical quantities. Guided by theoretical interest and pure curiosity, we ask how classical mechanics survives at the level of mean values while stochastic violations imprint themselves on fluctuations.
In this paper, we introduce a minimal model consisting of a two-body system with stochastically broken reciprocity, and then investigate its statistical signatures such as correlation functions, the evolution of probability distribution functions (PDFs), the energy gain and the entropy production, and so on. The remaining content of this paper is organized as follows. In Sec. II, we specifically describe a two-particle system with stochastically broken reciprocity, and derive the dynamic equations of motion for the system’s center of mass (COM) and the two-body relative position based on Newton’s second law. In Sec. III, we computes position-, velocity-, and cross-correlation functions for both modes of motion. In Sec. IV, we derive three Fokker-Planck equations governing the PDFs of the COM motion, the relative motion, and their joint evolution, and use them to track energy and entropy changes. In Sec. V, we immerse the system in a thermal bath and, under overdamped conditions, obtain the Smoluchowski equation for the relative position. In Sec. VI, we present an extremely minimal model in which the relative motion obeys classical mechanics exactly, while the stochastically broken reciprocity merely influence the fluctuating motion of the COM. The final section provides a brief summary and outlook.
@@
with $\delta_{\alpha\beta}$ the Kronecker symbol, $\delta(t-t^{\prime})$ the Dirac delta function, $\mathbf{I}$ the unit tensor. The superscript $T$ represents the transpose operation on vectors or matrices. Obviously, equations (2)–(4) ensure that the average relation (1) is satisfied.
@@
with $M\equiv m_{\mathrm{A}}+m_{\mathrm{B}}$ being the total mass of the system. Define the relative position vector between B and A
@@
Stochastic differential equations (10) and (11) combining with (14) and (15) govern the COM motion and the relative motion between B and A.
III correlation functions
The observable consequences of stochastically broken reciprocity are encoded in the correlation structure of the system. In what follows we examine the position-position, velocity-velocity, and position-velocity correlation functions for both the COM motion and the relative motion of the two particles. To be concrete, we restrict the discussion to one spatial dimension and take the interaction potential to be of harmonic form.
III.1 Correlation functions for the COM motion
@@
$X-X_{0}=V_{0}t+\int_{0}^{t}(t-\tau)\zeta_{1}(\tau)\mathrm{d}\tau,$ (24)
with $X_{0}$ the initial position of the COM. Considering Eq. (21), we immediately obtain $\left\langle V\right\rangle=\left\langle V_{0}\right\rangle$ and $\left\langle X-X_{0}\right\rangle=\left\langle V_{0}\right\rangle t$ from Eqs. (23) and (24). Hence, on average, the centre of mass executes uniform rectilinear motion.
We now calculate velocity autocorrelation function. Squaring Eq. (23) and considering Eq. (21), we have
@@
Repeating the similar procedure, we can achieve the velocity-displacement correlation
$\left\langle(X-X_{0})V\right\rangle=\left\langle V^{2}_{0}\right\rangle t+\frac{g_{11}}{2}t^{2}$ (27)
and the displacement (related to position) autocorrelation function
$\left\langle(X-X_{0})^{2}\right\rangle=\left\langle V^{2}_{0}\right\rangle t^{2}+\frac{g_{11}}{3}t^{3}.$ (28)
The linear, quadratic, and cubic time dependences in correlation functions (26)–(28) reflect the effect of stochastically broken reciprocity. When $g_{11}=0$, these results reduce to the familiar situation of uniform rectilinear motion in classical mechanics.
III.2 Correlation functions for the relative motion of B with respect to A
In one-dimensional situation, the relative position and velocity of B with respect to A are denoted as $x$ and $v$, respectively. According to Eqs. (11), (14) and (15), the equation of relative motion of B with respect to A can be expressed as
@@ III.2 Correlation functions for the relative motion of B with respect to A
Equations (29) and (33) describe a linear, damp-free harmonic oscillator driven by the stochastic force $\zeta_{2}(t)$. Following the standard procedure we achieve the solution
$\left(\begin{array}[]{c}v\\ \omega x\\ \end{array}\right)=U\left(\begin{array}[]{c}v_{0}\\ \omega x_{0}\\ \end{array}\right)+\int_{0}^{t}\zeta_{2}(\tau)\left(\begin{array}[]{c}\cos\omega(t-\tau)\\ \sin\omega(t-\tau)\\ \end{array}\right)\mathrm{d}\tau$ (34)
@@
We now evaluate the correlation functions. Define a covariance matrix
${\sigma}(t)\equiv\left(\begin{array}[]{cc}\langle v^{2}(t)\rangle&\omega\langle v(t)x(t)\rangle\\ \omega\langle x(t)v(t)\rangle&\omega\langle x^{2}(t)\rangle\\ \end{array}\right).$ (36)
Using Eqs. (31),(32), and the solution (34), one obtains
@@
The second term on the right-handed side of the above equation is the signature of stochastically broken reciprocity. When $g_{22}=0$, Eq. (37) returns to the familiar oscillator correlations in classical mechanics.
@@ IV.1 PDF for the COM motion
Deriving the evolution of the PDF from the stochastic equations of motion is a standard topic in modern statistical physics textbooks [43]. Following the method outlined in Ref. [43], we can straightforwardly derive the Fokker-Planck equation
@@
which describes the evolution of the PDF of the COM.
@@ IV.2 PDF for the two-body relative motion
which describes the evolution of the PDF corresponding to the two-body relative motion.
@@ IV.4 Energy gain
which implies that the system is experiencing an energy gain.
@@ IV.5 Entropy production
$\frac{\mathrm{d}S}{\mathrm{d}t}=\frac{1}{2}\int\left(\sum_{ij}g_{ij}s_{i}s_{j}\right)f\mathrm{d}^{4}\Gamma.$ (57)
@@ V Two-body system immersed in a thermal bath
with $D\equiv{T}/\nu+{{g_{\mathrm{A}}}}/{2\gamma_{\mathrm{A}}^{2}}+{{g_{\mathrm{B}}}}/{2\gamma_{\mathrm{B}}^{2}}$. According to the Einstein relation [43], we may define an effective temperature
@@
which is clearly larger than the bath temperature $T$.
Following the method in Ref. [43], we can easily derive the Smoluchowski equation corresponding to Eqs. (60)–(62). The PDF $P(\mathbf{r},t)$ for the two-body relative motion satisfies
@@
From the above equation, we observe that a steady state exists. The steady-state PDF has the Boltzmann form as
@@ VI Extremely minimal model
In the minimal model mention above, $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are assumed to be independent of each other. We may heuristically consider an extreme situation in which $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are not independent such that $\bm{\zeta}_{2}(t)=0$. In this sense, the two-body system is referred to as the extremely minimal model.
By setting Eq. (13) to be vanishing, we obtain a necessary condition: $\frac{\sqrt{g_{\mathrm{A}}}}{m_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=\frac{\sqrt{g_{\mathrm{B}}}}{m_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)$, which enlightens us to assume $\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=m_{\mathrm{A}}\sqrt{g}\bm{\xi}(t)$ and $\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)=m_{\mathrm{B}}\sqrt{g}\bm{\xi}(t)$ where $g$ is a positive constant quantity. To guarantee this assumption, we need to revisit the main equations in Sec. II. Equations (2) and (3) are rewritten as
@@
In the above discussion, we have proposed a minimal model consisting of a two-body system with stochastically broken reciprocity. Guided by Newton’s second law, we have derived two stochastic differential equations that govern the COM motion and the two-body relative motion. Based on these equations of motion, we have obtained the position-, velocity-, and cross-correlation functions for the COM motion and the two-body relative motion, respectively. Additionally, we have derived three Fokker-Planck equations, which respectively characterize the evolution of PDFs for the COM motion, the two-body relative motion, and the joint evolution of these two motions. Using these Fokker-Planck equations, we have analyzed the energy gain and entropy production behaviors in this two-body system arising from the stochastically broken reciprocity. We have further explored the two-body system in a constant-temperature thermal bath and determined the effective temperature of the system under the overdamped condition. The corresponding Smoluchowski equation, which describes the evolution of the PDF of the two-body relative position, yields the Boltzmann distribution with the effective temperature in the steady state. Finally, we have introduced an extremely minimal model where the relative motion adheres to the laws of classical mechanics, and the effect of stochastically broken reciprocity is solely manifested in the fluctuating motion of the COM.
@@
1) Experimental verification. A key question is whether the minimal model with stochastically broken reciprocity can be experimentally tested. As we have predicted, the correlation functions (Eqs. 26, 27, and 28) contain linear, quadratic, and cubic time-scaling terms, respectively. These scaling laws could be examined by observing the fluctuating motion of the COM in experiments. However, we regret that we cannot currently specify a real-world system where these effects might be observed, leaving this as a challenge for experimental physicists.
2) Extension to many-body systems. The framework of the present minimal model can be extended to many-body systems, where pairwise interactions are assumed to stochastically violate Newton’s third law following the same mechanism as in Eqs. 2 and 3. If no additional constraints are imposed on the entire system, the COM motion is expected to follow a rule similar to Eq. 10, meaning the correlation behaviors for the COM motion (discussed in Sec. IIIA) would remain valid. However, the motion of particles relative to the COM would be far more complex than in the two-body case, warranting further investigation.
@@ ACKNOWLEDGMENTS
The author is grateful to the support from the National Natural Science Foundation of China (Grant No. 12475032).

### v2

@@ Abstract
We introduce a minimal model consisting of a two-body system with stochastically broken reciprocity (i.e., random violation of Newton’s third law) and then investigate its statistical behaviors, including fluctuations of velocity and position, time evolution of probability distribution functions, energy gain, and entropy production. The effective temperature of this two-body system immersed in a thermal bath is also derived. Furthermore, we heuristically present an extremely minimal model where the relative motion adheres to the same rules as in classical mechanics, while the effect of stochastically broken reciprocity only manifests in the fluctuating motion of the center of mass.
Keywords: stochastically broken reciprocity, probability distribution function, energy gain, entropy production
@@ I Introduction
The strict breaking of reciprocity leads to odd viscosity or elasticity [25, 26, 29, 27, 28, 31, 30], unconventional phase transitions [32, 33, 34, 35, 36, 37, 38], and exotic transport behaviors [39, 40, 41, 42]. Rather than focusing on the aforementioned in-depth discussions on consequences of a strict violation of reciprocity, we aim to explore a more delicate scenario: Newton’s third law holds on average, but reciprocity is broken stochastically. Since the third law holds on average, we expect that the core conclusions in classical mechanics remain valid when expressed in terms of the mean values of relevant physical quantities. Stochastically broken reciprocity, however, plays a role in the fluctuations of these physical quantities. Guided by theoretical interest and pure curiosity, we ask how classical mechanics survives at the level of mean values while stochastic violations imprint themselves on fluctuations.
In this paper, we introduce a minimal model consisting of a two-body system with stochastically broken reciprocity, and then investigate its statistical signatures such as fluctuations of velocity and position, the evolution of probability distribution functions (PDFs), the energy gain and the entropy production, and so on. The remaining content of this paper is organized as follows. In Sec. II, we specifically describe a two-particle system with stochastically broken reciprocity, and derive the dynamic equations of motion for the system’s center of mass (COM) and the two-body relative position based on Newton’s second law. In Sec. III, we compute the mean square velocity and mean square displacement of the COM, and the covariance matrix of the two-body relative motion with a deterministic harmonic interaction. In Sec. IV, we derive three Fokker-Planck equations governing the PDFs of the COM motion, the relative motion, and their joint evolution, and use them to track the changes of energy and entropy. In Sec. V, we immerse the system in a thermal bath and, under overdamped conditions, obtain the Smoluchowski equation governing the PDF of the relative position. In Sec. VI, we present an extremely minimal model in which the relative motion obeys classical mechanics exactly, while the stochastically broken reciprocity merely influence the fluctuating motion of the COM. The final section provides a brief summary and outlook.
@@
with $\delta_{\alpha\beta}$ the Kronecker symbol, $\delta(t-t^{\prime})$ the Dirac delta function, $\mathbf{I}$ the unit tensor. The superscript “$\mathrm{T}$” represents the transpose operation on vectors or matrices. Obviously, equations (2)–(4) ensure that the average relation (1) is satisfied.
@@
with $M\equiv m_{\mathrm{A}}+m_{\mathrm{B}}$ being the total mass of the system. Define the relative position vector of B with respect to A
@@
Stochastic differential equations (10) and (11) combining with (14) and (15) govern the COM motion and the relative motion of B with respect A.
III Fluctuations of velocity and position
The observable consequences of stochastically broken reciprocity are encoded in the fluctuation behaviors of the system. In what follows we examine the mean square velocity and displacement for the COM motion, and the covariance matrix for the relative motion of the two particles. To be concrete, we restrict the discussion to one spatial dimension and take the interaction potential to be of harmonic form.
III.1 Mean square velocity and displacement for the COM motion
@@
$\Delta X\equiv X-X_{0}=V_{0}t+\int_{0}^{t}(t-\tau)\zeta_{1}(\tau)\mathrm{d}\tau,$ (24)
with $X_{0}$ the initial position of the COM. Considering Eq. (21), we immediately obtain $\left\langle V\right\rangle=\left\langle V_{0}\right\rangle$ and $\left\langle\Delta X\right\rangle=\left\langle V_{0}\right\rangle t$ from Eqs. (23) and (24). Hence, on average, the COM executes uniform rectilinear motion.
We now calculate the mean square velocity. Squaring Eq. (23) and considering Eq. (21), we have
@@
From this equation, we can directly write out the fluctuation of velocity, $\langle(V-\langle V\rangle)^{2}\rangle=\langle(V_{0}-\langle V_{0}\rangle)^{2}\rangle+g_{11}t$, which implies that the fluctuation of velocity increases linearly with time.
Repeating the similar procedure, we can achieve the mean square displacement:
$\left\langle(\Delta X)^{2}\right\rangle=\left\langle V^{2}_{0}\right\rangle t^{2}+\frac{g_{11}}{3}t^{3}.$ (27)
From the above equation, we can directly write out the fluctuation of displacement, $\langle(\Delta X-\langle\Delta X\rangle)^{2}\rangle=\langle(V_{0}-\langle V_{0}\rangle)^{2}\rangle t^{2}+({g_{11}}/{3})t^{3}$.
The linear and cubic time-dependent terms containing $g_{11}$ in Eqs. (26) and (27) reflect the effect of stochastically broken reciprocity. When $g_{11}=0$, these results reduce to the familiar situation of uniform rectilinear motion in classical mechanics.
III.2 Covariance matrix for the relative motion of B with respect to A
In one-dimensional situation, the relative position and velocity of B with respect to A are denoted as $x$ and $v$, respectively. According to Eqs. (11), (14) and (15), the equations of relative motion of B with respect to A can be expressed as
@@ III.2 Covariance matrix for the relative motion of B with respect to A
Equations (28) and (32) describe a linear, damp-free harmonic oscillator driven by the stochastic force $\zeta_{2}(t)$. Following the standard procedure, we achieve the solution:
$\left(\begin{array}[]{c}v\\ \omega x\\ \end{array}\right)=U\left(\begin{array}[]{c}v_{0}\\ \omega x_{0}\\ \end{array}\right)+\int_{0}^{t}\left(\begin{array}[]{c}\cos\omega(t-\tau)\\ \sin\omega(t-\tau)\\ \end{array}\right)\zeta_{2}(\tau)\mathrm{d}\tau$ (33)
@@
We now evaluate the fluctuation behaviors by introducing covariance matrix
${\sigma}(t)\equiv\left(\begin{array}[]{cc}\langle v^{2}\rangle-\langle v\rangle^{2}&\omega(\langle vx\rangle-\langle v\rangle\langle x\rangle)\\ \omega(\langle xv\rangle-\langle x\rangle\langle v\rangle)&\omega^{2}(\langle x^{2}\rangle-\langle x\rangle^{2})\\ \end{array}\right).$ (35)
Using Eqs. (30), (31), and the solution given by Eq. (33), we find that the covariance matrix evolves with time as follows:
@@
The second term on the right-handed side of the above equation is the signature of stochastically broken reciprocity. When $g_{22}=0$, Eq. (36) returns to the familiar evolution of the covariance matrix for oscillators in classical mechanics.
@@ IV.1 PDF for the COM motion
Deriving the evolution of the PDF from the stochastic equations of motion is a standard exercise in the textbook of modern statistical physics [43]. Following the method outlined in Ref. [43], we can straightforwardly derive the Fokker-Planck equation
@@
corresponding to Eqs. (19) and (20). This equation describes the evolution of the PDF of the COM.
@@ IV.2 PDF for the two-body relative motion
corresponding to Eqs. (28) and (29). This equation describes the evolution of the PDF of the two-body relative motion.
@@ IV.4 Energy gain
which implies that the system is undergoing an energy gain.
@@ IV.5 Entropy production
$\frac{\mathrm{d}S}{\mathrm{d}t}=\frac{1}{2}\int\left(\sum_{i=1}^{2}\sum_{j=1}^{2}g_{ij}s_{i}s_{j}\right)f\mathrm{d}^{4}\Gamma.$ (56)
@@ V Two-body system immersed in a thermal bath
It is noted that Eqs. (57) and (58), along with their underdamped counterparts, have been introduced by Kumar et al. [48] and Baule et al. [49] in the context of the Brownian inchworm model for self-propulsion. The exact solutions to these equations and comprehensive discussions on this model can be found in Refs. [48] and [49]. Herein, we outline only the key results regarding the two-body relative motion.
@@
with $D\equiv{T}/\nu+{{g_{\mathrm{A}}}}/{2\gamma_{\mathrm{A}}^{2}}+{{g_{\mathrm{B}}}}/{2\gamma_{\mathrm{B}}^{2}}$. Assuming that the Einstein relation [43] still holds, we may define an effective temperature
@@
which is clearly larger than the bath temperature $T$ since $\nu$, $g_{\mathrm{A}}$, $g_{\mathrm{B}}$ are positive quantities.
Following the method sketched in Ref. [43], we can readily derive the Smoluchowski equation corresponding to Eqs. (59)–(61). The PDF $P(\mathbf{r},t)$ for the two-body relative motion satisfies
@@
From the above equation, we observe that a steady state exists. Particularly, the steady-state PDF follows the Boltzmann distribution as
@@ VI Extremely minimal model
In the minimal model mention above, $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are assumed to be independent of each other. We may heuristically consider an extreme situation in which $\bm{\xi}_{\mathrm{A}}(t)$ and $\bm{\xi}_{\mathrm{B}}(t)$ are not independent such that $\bm{\zeta}_{2}(t)$ in Eq. (13) is vanishing. In this sense, the two-body system is referred to as the extremely minimal model.
By setting Eq. (13) to be vanishing, we obtain a necessary condition: $\frac{\sqrt{g_{\mathrm{A}}}}{m_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=\frac{\sqrt{g_{\mathrm{B}}}}{m_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)$, which enlightens us to assume $\sqrt{g_{\mathrm{A}}}\bm{\xi}_{\mathrm{A}}(t)=m_{\mathrm{A}}\sqrt{g}\bm{\xi}(t)$ and $\sqrt{g_{\mathrm{B}}}\bm{\xi}_{\mathrm{B}}(t)=m_{\mathrm{B}}\sqrt{g}\bm{\xi}(t)$ where $g$ is a positive constant quantity. To guarantee this assumption, we need to revisit the main equations in Sec. II. Specifically, equations (2) and (3) are rewritten as
@@
In the above discussion, we have proposed a minimal model consisting of a two-body system with stochastically broken reciprocity. Guided by Newton’s second law, we have derived two stochastic differential equations that govern the COM motion and the two-body relative motion, respectively. Based on these equations of motion, we have obtained the fluctuations of velocity and position for the COM motion and the two-body relative motion, respectively. Additionally, we have derived three Fokker-Planck equations, which respectively characterize the evolution of PDFs for the COM motion, the two-body relative motion, and the joint evolution of these two motions. Using these Fokker-Planck equations, we have analyzed the features of energy gain and entropy production in this two-body system arising from the stochastically broken reciprocity. We have further explored the two-body system in a constant-temperature thermal bath and determined the effective temperature of the system under the overdamped condition. The corresponding Smoluchowski equation, which describes the evolution of the PDF of the two-body relative position, yields the Boltzmann distribution with the effective temperature in the steady state. Finally, we have introduced an extremely minimal model where the relative motion adheres to the laws of classical mechanics, and the effect of stochastically broken reciprocity is solely manifested in the fluctuating motion of the COM.
@@
1) Experimental verification. A key question is whether the minimal model with stochastically broken reciprocity can be experimentally tested. As we have predicted, the mean square velocity (26) and the mean square displacement (27) contain linear and cubic time-scaling terms, respectively. These scaling laws could be examined by observing the fluctuating motion of the COM in experiments. However, we regret that we cannot currently specify a real-world system where these effects might be observed, leaving this as a challenge for experimental physicists.
2) Extension to many-body systems. The framework of the present minimal model can be extended to many-body systems, where pairwise interactions are assumed to stochastically violate Newton’s third law following the same mechanism as in Eqs. (2) and (3). If no additional constraints are imposed on the entire system, the COM motion is expected to follow a rule similar to Eq. (10), meaning the fluctuation behaviors for the COM motion (discussed in Sec. IIIA) would remain valid. However, the motion of particles relative to the COM would be far more complex than in the two-body case, warranting further investigation.
@@ ACKNOWLEDGMENTS
The author thanks Ananyo Maitra for drawing attention to references [48] and [49]. This work is supported by the National Natural Science Foundation of China (Grant No. 12475032).